# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v2

@@ Appendix A Graph Class Definitions
A graph $G$ is said to have genus $g$ if it can be embedded into an orientable surface of genus $g$, and it cannot be embedded into an orientable surface of smaller genus. Planar graphs are exactly those that have genus 00.
Up to homeomorphism, there is only one orientable surface of genus $g$, for any particular $g$. For $g=0$, it is the sphere. The orientable surface of genus $g>0$ can be obtained from the orientable surface of genus $g-1$ by attaching a “handle”. Alternatively, it can be expressed as $g$ “doughnuts” merged together. See Figure 6 for examples of these surfaces, and see Figure 7 for an example of a graph drawn on a genus-$3$ surface.
Figure 6: Examples of (left to right) a genus-$1$ surface, a genus-$2$ surface, a genus-$3$ surface, and a genus $10$ surface.
Figure 7: An example of a graph (left) and a drawing of the graph on a genus-$3$ surface (right).
@@ Appendix B Vertex Expansion is NP-hard
We now prove that $\frac{|\partial(S^{\prime\prime})|}{|S^{\prime\prime}|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}$: We can make all the modifications for vertices in $V_{1}$ “first”, and count the changes to the fraction, and then make the modifications for the vertices in $V_{2}$ and count the remaining changes. For the vertices $v^{i}\in V_{2}$ in $S^{\prime\prime}$ but not in $S^{\prime}$, they contributed at least $1$ to the numerator and 00 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 00 to the numerator and $1$ to the denominator. Henceforth, we may assume that $S^{\prime}$ does not “cut” individual sets $\{v_{1},\dots,v_{k}\}$. For the vertices $e=uv\in V_{2}$ not in $S^{\prime}$ but in $S^{\prime\prime}$, such vertices must have all their neighbours included in $S^{\prime\prime}$. Hence, they contributed at least $1$ to the numerator and 00 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 00 to the numerator and $1$ to the denominator. For the vertices $e=uv\in V_{2}$ in $S^{\prime}$ but not in $S^{\prime\prime}$, such vertices contributed $1$ to the denominator and $k$ to the numerator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute $1$ to the numerator and 00 to the denominator. We claim that each of these replacements (whose sequence eventually results in $S^{\prime\prime}$) decreases the value of the fraction. Let $a$ be the total size of $|\partial(S^{\prime})|$ excluding a clique of vertices in $V_{1}$, and let $b\geq 1$ be the total size of $|S^{\prime}|$ excluding a bisection vertex in $V_{2}$. We need to prove that $\frac{a+k}{b+1}\geq\frac{a+1}{b}$. If we assume $k\geq n^{2}+n+1$ (which is a polynomial in $n$) then, $a\leq n^{2}$ and $b\leq n$, so $k\geq n^{2}+n+1\implies kb\geq n^{2}+n+1\geq a+b+1\implies ab+kb\geq ab+a+b+1\implies(a+k)b\geq(a+1)(b+1)\implies\frac{a+k}{b+1}\geq\frac{a+1}{b}$.
@@ Appendix C Deferred Proof of Subdivision Lemma
Denote $G=(V,E)$, $n:=|V|$, $H:=G^{(k)}=(V^{\prime},E^{\prime})$, and $n^{\prime}:=|V^{\prime}|$. For the remainder of this proof, we will use the notation $\operatorname{\mathcal{O}}_{\Delta}()$ and $\Theta_{\Delta}()$ to hide polynomial factors in $\Delta$. Note that the maximum degree of $H$ is $\max\{6,\Delta\}$, since no new vertices of degree $>6$ are added from a hexagonal subdivision, nor does any existing vertex increase have its degree changed. Each edge in $G$ is split into $2^{k}$ pieces in $H$, and each triangle in $G$ is partitioned into $4^{k}$ triangles in $H$. The number of triangles incident to any vertex of degree $>6$ remains constant during subdivision. Hence, $\frac{n^{\prime}}{n}\in\Theta_{\Delta}(4^{k})$. Let $y_{H},f_{H}$ denote the optimal solution to $\gamma^{(1)}(H)$. Assume without loss of generality that $\sum_{x^{\prime}\in V^{\prime}}|f_{H}|^{2}=1$. This assumption will allow us to slightly simplify some later steps.

### v3

@@ Appendix A Graph Class Definitions
A graph $G$ is said to have genus $g$ if it can be embedded into an orientable surface of genus $g$, and it cannot be embedded into an orientable surface of smaller genus. Planar graphs are exactly those that have genus 0.
Up to homeomorphism, there is only one orientable surface of genus $g$, for any particular $g$. For $g=0$, it is the sphere. The orientable surface of genus $g>0$ can be obtained from the orientable surface of genus $g-1$ by attaching a “handle”. Alternatively, it can be expressed as $g$ “doughnuts” merged together. See Figure 6 for examples of these surfaces, and see Figure 7 for an example of a graph drawn on a genus-333 surface.
Figure 6: Examples of (left to right) a genus-111 surface, a genus-222 surface, a genus-333 surface, and a genus 101010 surface.
Figure 7: An example of a graph (left) and a drawing of the graph on a genus-333 surface (right).
@@ Appendix B Vertex Expansion is NP-hard
We now prove that $\frac{|\partial(S^{\prime\prime})|}{|S^{\prime\prime}|}\leq\frac{|\partial(S^{\prime})|}{|S^{\prime}|}$: We can make all the modifications for vertices in $V_{1}$ “first”, and count the changes to the fraction, and then make the modifications for the vertices in $V_{2}$ and count the remaining changes. For the vertices $v^{i}\in V_{2}$ in $S^{\prime\prime}$ but not in $S^{\prime}$, they contributed at least 111 to the numerator and 0 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 0 to the numerator and 111 to the denominator. Henceforth, we may assume that $S^{\prime}$ does not “cut” individual sets $\{v_{1},\dots,v_{k}\}$. For the vertices $e=uv\in V_{2}$ not in $S^{\prime}$ but in $S^{\prime\prime}$, such vertices must have all their neighbours included in $S^{\prime\prime}$. Hence, they contributed at least 111 to the numerator and 0 to the denominator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 0 to the numerator and 111 to the denominator. For the vertices $e=uv\in V_{2}$ in $S^{\prime}$ but not in $S^{\prime\prime}$, such vertices contributed 111 to the denominator and $k$ to the numerator under $S^{\prime}$, and under $S^{\prime\prime}$ they contribute 111 to the numerator and 0 to the denominator. We claim that each of these replacements (whose sequence eventually results in $S^{\prime\prime}$) decreases the value of the fraction. Let $a$ be the total size of $|\partial(S^{\prime})|$ excluding a clique of vertices in $V_{1}$, and let $b\geq 1$ be the total size of $|S^{\prime}|$ excluding a bisection vertex in $V_{2}$. We need to prove that $\frac{a+k}{b+1}\geq\frac{a+1}{b}$. If we assume $k\geq n^{2}+n+1$ (which is a polynomial in $n$) then, $a\leq n^{2}$ and $b\leq n$, so $k\geq n^{2}+n+1\implies kb\geq n^{2}+n+1\geq a+b+1\implies ab+kb\geq ab+a+b+1\implies(a+k)b\geq(a+1)(b+1)\implies\frac{a+k}{b+1}\geq\frac{a+1}{b}$.
@@ Appendix C Deferred Proof of Subdivision Lemma
Denote $G=(V,E)$, $n:=|V|$, $H:=G^{(k)}=(V^{\prime},E^{\prime})$, and $n^{\prime}:=|V^{\prime}|$. For the remainder of this proof, we will use the notation $\operatorname{\mathcal{O}}_{\Delta}()$ and $\Theta_{\Delta}()$ to hide polynomial factors in $\Delta$. Note that the maximum degree of $H$ is $\max\{6,\Delta\}$, since no new vertices of degree >6>6> 6 are added from a hexagonal subdivision, nor does any existing vertex increase have its degree changed. Each edge in $G$ is split into $2^{k}$ pieces in $H$, and each triangle in $G$ is partitioned into $4^{k}$ triangles in $H$. The number of triangles incident to any vertex of degree >6>6> 6 remains constant during subdivision. Hence, $\frac{n^{\prime}}{n}\in\Theta_{\Delta}(4^{k})$. Let $y_{H},f_{H}$ denote the optimal solution to $\gamma^{(1)}(H)$. Assume without loss of generality that $\sum_{x^{\prime}\in V^{\prime}}|f_{H}|^{2}=1$. This assumption will allow us to slightly simplify some later steps.